. If the stress produced in stretching a wire is 5.00 x 10 6 N/m2 by an applied force of

10.0 N, what is the cross-sectional area of the wire?  


1
Expert's answer
2022-05-20T04:53:53-0400
"\sigma=\frac{F}{A}"

"A=\frac{F}{\sigma}=\frac{10.0\;\rm N}{5.00*10^6\:\rm N/m^2}\\
=2*10^{-6}\:\rm m^2=2\: mm^2"